Sandy Bank, Inc., makes one model of wooden canoe. Partial information is given below.

Required:

1. Complete the following table.

2. Suppose Sandy Bank sells its canoes for $520 each. Calculate the contribution margin per canoe and the contribution margin ratio.

3. This year Sandy Bank expects to sell 810 canoes. Prepare a contribution margin income statement for the company.

4. Calculate Sandy Bank’s break-even point in units and in sales dollars.

5. Suppose Sandy Bank wants to earn $70,000 profit this year. Calculate the number of canoes that must be sold to achieve this target.

Complete the following table. (Round your "Cost per Unit" answers to 2 decimal places.)

Number of Canoes Produced and Sold 480 600 760
Total costs
Variable Costs $76,320
Fixed Costs 155,520
Total Costs $231,840
Cost per Unit
Variable Cost per Unit
Fixed Cost per Unit
Total Cost per Unit
Answer #1
Total fixed costs and variable cost per unit remains constant within the relevant range

Number of Canoes Produced and Sold

480 600 760
Total costs
Variable Costs     76,320.00     95,400.00     120,840.00
Fixed Costs 155,520.00 155,520.00     155,520.00
Total Costs 231,840.00 250,920.00     276,360.00
Cost per Unit
Variable Cost per Unit          159.00          159.00            159.00
Fixed Cost per Unit          324.00          259.20            204.63
Total Cost per Unit          483.00          418.20            363.63
1.Contribution margin per unit = Selling price per unit - Variable cost per unit
=520-159 = $361
CM Ratio = CM/Sales
69.42%
Contribution Margin Income Statement
Sales Revenue 421200
Less: Variable costs 128,790.00
Contribution Margin 292,410.00
Fixed costs 155,520.00
Operating Income 136,890.00
Break even point in units = Fixed costs/CM Per unit
                                                         430.80
431 canoes
Sales Dollars = Fixed costs/CM Ratio
                                                  224,017.73
3.Number of canoes = (70,000+155520)/361
624.7091413 canoes